OpenStudy (anonymous):

a cube has side length 15 kilometers. What is its surface area?

OpenStudy (cggurumanjunath):

surface area of cube = 6*a*a in km^2 where a=15km

OpenStudy (mathmale):

Yes: if the length of one side is a = 15 km, the surface area of this cube is A=6*a^2, or A=6(15 km)^2.

OpenStudy (anonymous):

A cube has 6 sides. The surface area of one side would be 15 * 15 since all the side lengths are the same on a cube. 15 * 15 = 225 times the 6 sides = 1,350 km²

OpenStudy (anonymous):

thank you guys!!!

OpenStudy (anonymous):

You're welcome!! Xo!